# SAFEx: Analyzing Vulnerabilities of MoE-Based LLMs via Stable Safety-critical Expert Identification

### v1

@@ Abstract
Large language models based on Mixture-of-Experts have achieved substantial gains in efficiency and scalability, yet their architectural uniqueness introduces underexplored safety alignment challenges. Existing safety alignment strategies, predominantly designed for dense models, are ill-suited to address MoE-specific vulnerabilities. In this work, we formalize and systematically study MoE model’s positional vulnerability—the phenomenon where safety-aligned behaviors rely on specific expert modules, revealing critical risks inherent to MoE architectures. To this end, we present SAFEx, an analytical framework that robustly identifies, characterizes, and validates the safety-critical experts using a novel Stability-based Expert Selection (SES) algorithm. Notably, our approach enables the explicit decomposition of safety-critical experts into distinct functional groups, including those responsible for harmful content detection and those controlling safe response generation. Extensive experiments on mainstream MoE models, such as recently released Qwen3-MoE demonstrated that their intrinsic safety mechanisms heavily rely on a small subset of positional experts. Disabling these experts significantly compromised the models’ ability to refuse harmful requests. For Qwen3-MoE with $6144$ experts (in FNN layer), we find that disabling as few as 12 identified safety-critical experts can cause the refusal rate to drop by $22\%$, demonstrating the disproportionate impact of a small set of experts on overall model safety.
footnotetext: Co-corresponding authors: lijq@szu.edu.cn, wubingzheagent@gmail.comfootnotetext: These authors contributed equally to this work.
1 School of Artificial Intelligence, Shenzhen University
2 National Engineering Laboratory for Big Data System Computing Technology, Shenzhen University
3 College of Computer Science and Software Engineering, Shenzhen University
4 ByteDance Inc
@@ 1 Introduction
Large language models (LLMs) based on Mixture-of-Experts (MoE) architectures, such as Mixtral [1], DeepSeek-R1 [2], and Qwen3-MoE [3], have achieved remarkable advances on a wide range of complex tasks, significantly improving efficiency and scalability by routing inputs dynamically across multiple specialized expert modules. Despite these successes, the distinctive architectural characteristics of MoE-based models raise unique and underexplored safety issues.
Recent works on safety research of MoE-based LLMs have started to emerge [4, 5], yet remain nascent and primarily focus on exploiting MoE-specific architectural vulnerabilities to attack LLM models. For example, BadMoE [6] identifies and exploits rarely activated experts, referred to as "dormant experts", to successfully execute effective adversarial attacks, highlighting significant safety vulnerabilities inherent to MoE-based architectures. However, these existing studies predominantly concentrate on demonstrating potential vulnerabilities to specific attacks and leave a critical gap in comprehensively understanding how existing safety alignment mechanisms influence the internal behaviors of MoE-based LLMs.
Figure 1 intuitively illustrates that the intrinsic safety-aligned behaviors of MoE-based LLMs strongly depend on specific positional experts, a phenomenon we define as positional vulnerability. Specifically, as Figure 1 shows: when an MoE model processes a typical harmful input, certain experts are consistently activated at a high frequency (represented as bar charts) during decoding (highlighted in red Figure 1 (a)); Notably, when we intentionally freeze (inactivate) these frequently activated experts during inference (illustrated in gray), the model immediately begins generating unsafe responses (Figure 1 (b)); Furthermore, applying advanced jailbreak methods [7, 8, 9] to the same harmful input triggers a dramatic shift in the expert activation distribution, resulting in the activation of entirely different experts and subsequently producing unsafe outputs (Figure 1 (c)).
@@
    To enhance the reliability of expert activation probability estimation, we design a stability-based statistical selection algorithm inspired by Stability-based Feature Selection (SES). This approach provides robust identification and validation of safety-critical positional experts, significantly improving the interpretability and reliability of analytical outcomes.
@@ 2.1 Dataset Construction
To systematically analyze the safety-alignment behaviors of MoE-based LLMs, we carefully construct a specialized evaluation dataset. Our dataset consists of three distinct groups designed to cover different input scenarios relevant to model safety alignment:
  • Regular Group ($\mathcal{D}_{Regular}$): This group consists of original harmful requests. To ensure comprehensive coverage, we uniformly sample harmful prompts from multiple predefined harmful categories (e.g., fraud, health consultation, illegal activities) from existing different benchmark datasets [11, 12, 13, 14, 15]. The detailed distribution of harmful content categories in $\mathcal{D}_{Regular}$ and their corresponding data sources are illustrated in the appendix.
  • Jailbreak Group ($\mathcal{D}_{Jailbreak}$): Corresponding directly to the Regular group, the Jailbreak group contains harmful requests transformed using advanced jailbreak techniques—such as semantic paraphrasing [16], adversarial perturbations, or context reframing—to bypass the model’s safety mechanisms. Each prompt in this group is derived from the Regular group, enabling direct comparative analyses.
@@ 2.2 Safety-related Expert Activation Probability Modeling
The core task in our proposed workflow is to model the conditional probability distribution of expert activation states given specific input prompt types (e.g., harmful requests) as shown in Figure 2 (a). Formally, for an expert $e$, given a prompt distribution $\mathcal{X}$ sharing the same attribute (e.g., harmful content), we aim to model the conditional activation probability $p(e\mid\mathcal{X})$, as the probability of activating the expert $e$ during inference over a dataset.
The most straightforward way to estimate this conditional distribution is through empirical frequency counting on a specific prompt dataset $X$ sampling from $\mathcal{X}$, and computing the empirical estimation $p(e|X)$. Specifically, given a particular MoE architecture and a set of prompts $X=\{x^{(1)},x^{(2)},\dots,x^{(N)}\}$, we perform inference and record expert activation states at each token position. Given an attention block indexed by $l$, each token position $t$ within the full inference process (including encoding and decoding phases) activates exactly one expert from the set of experts =$\mathcal{E}=\{e_{1},e_{2},\dots,e_{M}\}$. We denote the activated expert at layer $l$, token position $t$, for prompt $x^{(i)}$ as $e_{l,t}^{(i)}$. Thus, the conditional probability of expert $e$ activation can be empirically estimated as:
$p(e\mid X)=\frac{\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}\mathbb{I}(e_{l,t}^{(i)}=e)}{\sum_{i=1}^{N}T_{i}}$
where $\mathbb{I}(\cdot)$ is the indicator function, $T_{i}$ denotes the total number of tokens generated in response to prompt $x^{(i)}$, and $N$ is the total number of prompts in the dataset $X$.
Stability-based Expert Selection.  However, the direct frequency counting method inherently depends on specific empirical input distributions, potentially limiting the generalizability of our findings. To alleviate this dependency and achieve a more robust estimation, we propose an alternative sampling strategy motivated by stability selection, a principled statistical approach [10]. Specifically, given an underlying input distribution $\mathcal{X}$, we independently draw multiple empirical datasets $\{X^{(1)},X^{(2)},\dots,X^{(K)}\}$, each consisting of samples independently and identically distributed (i.i.d.) from $\mathcal{X}$. In practical implementation, we propose to sample without replacement from a sufficiently large dataset (e.g., $\mathcal{D}_{Regular}$). For each empirical dataset $X^{(k)}$, we independently estimate the conditional probability distribution of expert activation as follows:
$p(e\mid X^{(k)})=\frac{\sum_{x^{(i)}\in X^{(k)}}\sum_{t=1}^{T_{i}}\mathbb{I}(e_{l,t}^{(i)}=e)}{\sum_{x^{(i)}\in X^{(k)}}T_{i}}$
where $e_{l,t}^{(i)}$ denotes the expert activated at decoding step $t$ for input $x^{(i)}$, and $T_{i}$ is the length of the decoded sequence for input $x^{(i)}$.
Next, we aggregate these dataset-level estimates by identifying the intersection of top-ranked experts across all independently sampled empirical datasets, thereby obtaining a stable set of frequently activated ("head") experts robust to variations in the empirical input distribution. Formally, we define the set of top experts from each empirical dataset $X^{(k)}$ as:
$\mathcal{E}_{top}^{(k)}(X)=\text{Top-N}_{e}\left(p(e\mid X^{(k)})\right),$
[…]
$\mathcal{E}_{top}(X)=\bigcap_{k=1}^{K}\mathcal{E}_{top}^{(k)}(X).$
This revised procedure ensures that our identified expert set is less sensitive to specific dataset biases and more accurately reflects robust expert activation patterns inherent to the underlying data distribution $\mathcal{X}$.
We apply the above modeling procedure separately to each of the three constructed data groups—namely, the Regular group $\mathcal{D}_{Regular}$, the Jailbreak group $\mathcal{D}_{Jailbreak}$, and the Benign group $\mathcal{D}_{Benign}$. We derive three distinct expert activation sets: $\mathcal{E}_{top}(\mathcal{D}_{Regular})$, $\mathcal{E}_{top}(\mathcal{D}_{Jailbreak})$, and $\mathcal{E}_{top}(\mathcal{D}_{Benign})$, which represent experts activated consistently within Regular, Jailbreak, and Benign scenarios, respectively.
@@
Expert Activation Visualization and Analysis.  We apply the above expert activation modeling framework systematically across three representative MoE-based LLMs: Mixtral-8x7B-Instruct-v0.1 [1], deepseek-moe-16b-chat [17], and Qwen3-30B-A3B [3]. For each model, we identify and visualize the layer and expert indices of the most frequently activated experts, together with their activation probabilities, as shown in Figure 3. Each plot highlights distinctive activation patterns across the three evaluation datasets, providing intuitive insights into the functional specialization of expert modules.
@@
Figure 3: Activation probability visualization of $\mathcal{E}_{\mathrm{top}}(\mathcal{D}_{\mathrm{Regular}})$, $\mathcal{E}_{\mathrm{top}}(\mathcal{D}_{\mathrm{Jailbreak}})$, and $\mathcal{E}_{\mathrm{top}}(\mathcal{D}_{\mathrm{Benign}})$ for three MoE models. Dashed lines denote the theoretical mean activation probability under each MoE configuration.
@@
Figure 3 shows the activation probability distributions of top-ranked experts across different MoE models and their corresponding datasets (this figure only shows 3 models due to page limits, more results can be found in Appendix). The horizontal dashed lines represent the theoretical mean activation probabilities under each MoE configuration. For the Regular dataset (first column), expert activation patterns differ significantly across architectures. Both Qwen3-30B-A3B and Mixtral-8x7B-Instruct-v0.1 exhibit prominent head-expert concentration—certain experts activate at substantially higher probabilities than the model average—while deepseek-moe-16b-chat shows an even stronger concentration with larger variance across experts. In contrast, Mixtral’s routing appears more balanced, with activation levels closely aligning with the mean. This discrepancy likely reflects differences in post-training safety alignment strategies among the models.
@@ 2.3 Expert Functional Categorization and Localization
Building upon the expert activation statistics obtained previously, we further employ straightforward set-based analyses to precisely identify and categorize expert modules according to their specific functional roles in safety alignment, as shown in Figure 2 (b). Specifically, we define two key functional expert groups:
@@
    $\mathcal{E}_{\text{id}}=\mathcal{E}_{top}(\mathcal{D}_{Regular})\cap\mathcal{E}_{top}(\mathcal{D}_{Jailbreak})$
@@
    $\mathcal{E}_{\text{ctrl}}=\mathcal{E}_{top}(\mathcal{D}_{Regular})-\mathcal{E}_{top}(\mathcal{D}_{Jailbreak})$
@@ 2.4 Linear Probing Experiment for $\mathcal{E}_{\text{id}}$
$\hat{y}=\sigma(\mathbf{W}\times\mathbf{h}_{\text{id}}(x)+b),\quad\hat{y}\in\{0,1\}$
where $\sigma$ denotes a logistic sigmoid function, $\mathbf{W},b$ are the classifier parameters. In practice, we use the average (across all input tokens) output from FFN layer of the selected expert to construct $\mathbf{h}_{\mathrm{id}}(x)$.
@@
Finally, we present these linear probes performance metrics across different models using box plots in Figure 4. As shown in the figure, linear probes constructed from experts within $\mathcal{E}_{\text{id}}$ consistently outperform the randomly selected experts ($\mathcal{E}_{\text{random}}$) across various metrics and model architectures. This substantial performance gap quantitatively confirms the distinctive role and discriminative capacity of the identified expert group ($\mathcal{E}_{\text{id}}$) in recognizing and distinguishing harmful content within MoE-based LLMs.

### v2

@@ Abstract
Large language models with Mixture-of-Experts (MoE) architectures achieve efficiency and scalability, yet their routing mechanisms introduce safety alignment challenges insufficiently addressed by techniques developed for dense models. In this work, the MoE-specific safety risk of positional vulnerability—that safety-aligned behaviors rely on specific expert modules—is formalized and systematically analyzed. An analytical framework, SAFEx, is presented to robustly identify, characterize, and validate safety-critical experts via a stability-based expert selection procedure, and to decompose them into two functional groups: the Harmful Content Detection Group (HCDG), which specializes in identifying and recognizing harmful content within user inputs, and the Harmful Response Control Group (HRCG), which specializes in controlling and enforcing model behaviors to generate appropriate safety responses. Expert-level interventions are conducted to probe causality and to test mitigation. Targeted masking of SAFEx-selected experts reveals that safety behavior is highly concentrated. On Qwen3-30B-A3B, configured with 48 MoE-FFN layers and 128 experts per layer under top-8 routing ($48\times 128=6{,}144$ experts in total), disabling 12 selected experts reduces the refusal rate by $22\%$. In addition, lightweight adaptation is performed using LoRA under three configurations—the HRCG, the union of HCDG and HRCG, and all experts—and the resulting updates are composed through negative weight merging targeted at the HRCG, leading to improved refusal under adversarial prompts without full-model retraining. These results establish positional vulnerability as a distinct MoE-specific safety challenge and provide a practical, compute-efficient pathway for expert-level safety interventions within routed architectures.
@@ 1 Introduction
Large language models (LLMs) using Mixture-of-Experts (MoE) architectures, such as Mixtral [1], DeepSeek-R1 [2], and Qwen3-MoE [3], have achieved significant progress on complex tasks by routing inputs across specialized experts, greatly improving efficiency and scalability. However, the particular architecture of MoE-based models raises unique and underexplored safety issues.
Recent works on safety research of MoE-based LLMs have started to emerge [4, 5], yet remain nascent and primarily focus on exploiting MoE-specific architectural vulnerabilities to attack LLM models. For example, BadMoE [6] identifies and exploits rarely activated experts, referred to as “dormant experts”, to successfully execute effective adversarial attacks, highlighting significant safety vulnerabilities inherent to MoE-based architectures. However, these existing studies predominantly concentrate on demonstrating potential vulnerabilities to specific attacks and leave a critical gap in comprehensively understanding how existing safety alignment mechanisms influence the internal behaviors of MoE-based LLMs.
Figure 1 intuitively illustrates that the intrinsic safety-aligned behaviors of MoE-based LLMs strongly depend on specific positional experts, a phenomenon we define as positional vulnerability. Specifically, as Figure 1 shows: when an MoE model processes a typical harmful input, certain experts are consistently activated at a high frequency (represented as bar charts) during decoding (highlighted in red Figure 1 (a)); Notably, when we intentionally inactivate these frequently activated experts during inference (illustrated in gray), the model immediately begins generating unsafe responses (Figure 1 (b)); Furthermore, applying advanced jailbreak methods [7, 8, 9] to the same harmful input triggers a dramatic shift in the expert activation distribution, resulting in the activation of entirely different experts and subsequently producing unsafe outputs (Figure 1 (c)).
@@
    To enhance the reliability of expert activation probability estimation, we design a Stability-based Expert Selection (SES) algorithm inspired by Stability-based Feature Selection. This procedure robustly identifies and validates safety-critical positional experts, thereby improving the interpretability and reliability of our analysis.
@@ 2.1 Dataset Construction
To systematically analyze the safety-alignment behaviors of MoE-based LLMs, we carefully construct a specialized evaluation dataset. Our dataset is organized into three distinct groups, with 500 samples per group, designed to capture a diverse range of input scenarios relevant to model safety alignment:
  • Regular Group ($\mathcal{D}_{Regular}$): This group consists of original harmful requests. To ensure comprehensive coverage, we uniformly sample harmful prompts from multiple predefined harmful categories (e.g., fraud, health consultation, illegal activities) from existing different benchmark datasets [11, 12, 13, 14, 15]. The detailed distribution of harmful content categories in $\mathcal{D}_{Regular}$ and their corresponding data sources are illustrated in the Appendix Figure 6.
  • Jailbreak Group ($\mathcal{D}_{Jailbreak}$): Derived from the Regular group, this set contains harmful requests rewritten with three jailbreak strategies (semantic paraphrasing [16], adversarial perturbations, and context reframing) to evade the model’s safety mechanisms. All adversarial prompts were generated by an external model (DeepSeek-V3/R1) in a black box setting without access to routing paths or gradients, allowing direct comparison with the Regular group.
@@ 2.2 Safety-related Expert Activation Probability Modeling
Notation. Consider an MoE model with $L$ MoE-FFN layers indexed by $l\in\{1,\dots,L\}$. For each layer, the number of experts is fixed at $K$; we denote the layer-$l$ expert set as $\mathcal{E}_{l}=\{e_{l,1},\dots,e_{l,K}\}$ with $|\mathcal{E}_{l}|=K$. The router selects exactly $k$ experts per token per layer (Top-$k$). Model-specific $(K,k)$ configurations are summarized in Appendix Table 3. Given a prompt distribution $\mathcal{X}$ and a dataset $X=\{x^{(1)},\dots,x^{(N)}\}\sim\mathcal{X}$, let $T_{i}$ denote the number of processed tokens for $x^{(i)}$. At token position $t$ of $x^{(i)}$ and layer $l$, let $A^{(i)}_{l,t}\subseteq\mathcal{E}_{l}$ be the Top-$k$ set with $|A^{(i)}_{l,t}|=k$.
Objective. For an expert $e\in\mathcal{E}_{l}$, we aim to model the conditional activation probability $p_{l}(e\mid\mathcal{X})$ under prompts sharing a given attribute (e.g., harmful requests), as illustrated in Figure 2(a).
Estimator. A simple and unbiased estimator is empirical frequency counting over $X$:
$p_{l}(e\mid X)\;=\;\frac{\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}\mathbb{I}\!\bigl(e\in A^{(i)}_{l,t}\bigr)}{\sum_{i=1}^{N}T_{i}},\qquad\sum_{e\in\mathcal{E}_{l}}p_{l}(e\mid X)=k,$
where $\mathbb{I}(\cdot)$ is the indicator function. Under perfectly uniform routing, the theoretical mean activation equals $k/K$. For example, for Qwen3-30B-A3B with $k=8$ and $K=128$, the baseline is $8/128=0.0625$. When a single scalar summary is needed, we report $p_{l}(e\mid X)$ for representative layers or the layer-averaged
$\frac{1}{L}\sum_{l=1}^{L}p_{l}(e\mid X)$
across all $L$ MoE layers.
Stability-based Expert Selection.
Direct frequency counting may overfit a particular empirical sample $X$. To improve robustness, we adopt a stability-selection-style resampling strategy [10].
Resampling protocol. Draw $S$ bootstrap subsets $\{X^{(1)},\dots,X^{(S)}\}$ from a sufficiently large pool (sampling with replacement). For large pools, sampling without replacement approximates independence in practice.
Per-resample estimation. For each $X^{(s)}$ we estimate
$p_{l}\!\left(e\mid X^{(s)}\right)\;=\;\frac{\sum_{x^{(i)}\in X^{(s)}}\sum_{t=1}^{T_{i}}\,\mathbb{I}\!\bigl(e\in A^{(i)}_{l,t}\bigr)}{\sum_{x^{(i)}\in X^{(s)}}T_{i}},\qquad\sum_{e\in\mathcal{E}_{l}}p_{l}(e\mid X^{(s)})=k.$
Stable top-$N_{e}$ experts. Let $\mathrm{Top}\text{-}N_{e}(\cdot)$ return the top-$N_{e}$ experts ranked by $p_{l}(\cdot)$ (per-layer, or using a layer-averaged score). Unless noted, we set $N_{e}=\lfloor\alpha K\rfloor$ for a fixed $\alpha\in(0,1]$. Define
$\mathcal{E}_{\mathrm{top}}^{(s)}=\mathrm{Top}\text{-}N_{e}\!\bigl(p_{l}(e\mid X^{(s)})\bigr),\qquad\mathcal{E}_{\mathrm{top}}=\bigcap_{s=1}^{S}\mathcal{E}_{\mathrm{top}}^{(s)}.$
This intersection yields a set of “stable” (frequently activated) experts that is less sensitive to dataset-specific fluctuations and better reflects activation patterns inherent to $\mathcal{X}$.
Group-wise application. We apply the procedure separately to the Regular, Jailbreak, and Benign groups, i.e., $\mathcal{D}_{Regular}$, $\mathcal{D}_{Jailbreak}$, and $\mathcal{D}_{Benign}$, obtaining three expert sets: $\mathcal{E}_{\mathrm{top}}(\mathcal{D}_{Regular})$, $\mathcal{E}_{\mathrm{top}}(\mathcal{D}_{Jailbreak})$, and $\mathcal{E}_{\mathrm{top}}(\mathcal{D}_{Benign})$. These serve as the basis for downstream analyses linking expert functionality to harmful-content identification and safety-aligned response control.
@@
Expert Activation Visualization and Analysis.  We apply the above expert activation modeling framework systematically across three representative MoE‐based LLMs: Mixtral-8x7B-Instruct-v0.1 [1], deepseek-moe-16b-chat [17], and Qwen3-30B-A3B [3]. For each model, we identify and visualize the layer and expert indices of the most frequently activated experts, together with their activation probabilities, as shown in Figure 3. Each plot highlights distinctive activation patterns across the three evaluation datasets, providing intuitive insights into the functional specialization of expert modules.
@@
Figure 3: Activation probability visualization of $\mathcal{E}_{\mathrm{top}}(\mathcal{D}_{\mathrm{Regular}})$, $\mathcal{E}_{\mathrm{top}}(\mathcal{D}_{\mathrm{Jailbreak}})$, and $\mathcal{E}_{\mathrm{top}}(\mathcal{D}_{\mathrm{Benign}})$ for three MoE models. The horizontal dashed lines represent the theoretical mean activation probability $k/K$ under each MoE configuration.
@@
Figure 3 presents the activation probability distributions of top-ranked experts across three representative MoE models (additional results are provided in the Appendix). Although the curves appear smooth, the conservation of total activations within each layer implies that head-expert elevations above the theoretical mean activation probability inevitably suppress a long tail of experts, revealing non-uniform routing under jailbreak. For the Regular dataset (first column), expert activation patterns differ significantly across architectures. Both Qwen3-30B-A3B and Mixtral-8x7B-Instruct-v0.1 exhibit prominent head-expert concentration—certain experts activate at substantially higher probabilities than the model average—while deepseek-moe-16b-chat shows an even stronger concentration with larger variance across experts. In contrast, Mixtral’s routing appears more balanced, with activation levels closely aligning with the mean. This discrepancy likely reflects differences in post-training safety alignment strategies among the models.
@@ 2.3 Expert Functional Categorization and Localization
Building upon the expert activation statistics obtained previously, we further employ straightforward set-based analyses to precisely identify and categorize expert modules according to their specific functional roles in safety alignment, as shown in Figure 2 (b). Consistent with the abstract, we refer to the Harmful Content Detection Group (HCDG) and the Harmful Response Control Group (HRCG) as $\mathcal{E}_{\text{id}}$ and $\mathcal{E}_{\text{ctrl}}$, respectively, throughout the paper:
@@
    $\mathcal{E}_{\text{id}}=\mathcal{E}_{top}(\mathcal{D}_{Regular})\cap\mathcal{E}_{top}(\mathcal{D}_{Jailbreak}).$
@@
    $\mathcal{E}_{\text{ctrl}}=\mathcal{E}_{top}(\mathcal{D}_{Regular})-\mathcal{E}_{top}(\mathcal{D}_{Jailbreak}).$
@@ 2.4 Linear Probing Experiment for $\mathcal{E}_{\text{id}}$
[…]
where $\sigma$ denotes a logistic sigmoid function, and $\mathbf{W},b$ are the classifier parameters. In practice, we use the average (across all input tokens) output from the FFN layer of the selected expert to construct $\mathbf{h}_{\mathrm{id}}(x)$.
@@
Hyper-parameters for the linear probes. Each probe is an $L_{2}$-regularized logistic regression classifier trained with regularization strength $C=1.0$ using the lbfgs solver.
@@
Finally, we present these linear probes’ performance metrics across different models using box plots in Figure 4. As shown in the figure, linear probes constructed from experts within $\mathcal{E}_{\text{id}}$ consistently outperform the randomly selected experts ($\mathcal{E}_{\text{random}}$) across various metrics and model architectures. This substantial performance gap quantitatively confirms the distinctive role and discriminative capacity of the identified expert group ($\mathcal{E}_{\text{id}}$) in recognizing and distinguishing harmful content within MoE-based LLMs.